11)Solve the PDE (x+y
1
Expert's answer
2021-09-22T00:25:20-0400

Auxiliary equations are



"{dx\over x(y^2+z)}={dy \over -y(x^2+z)}={dz \over z(x^2-y^2)}"

By Choosing multipliers "x, y, -1," we get



"{xdx+ydy-dz\over x^2y^2+x^2z-x^2y^2-y^2z-x^2z+y^2z}={xdx+ydy-dz\over 0}"

Then



"x^2+y^2-2z=C_1"

By Choosing multipliers "1\/x, 1\/y, 1\/z," we get



"{{dx \over x}+{dx \over x}+{dz \over z}\over y^2+z-x^2-z+x^2-y^2}={{dx \over x}+{dx \over x}+{dz \over z}\over 0}"

Then



"\ln(xyz)=\ln (C_2)"

Or



"xyz=C_2"

Parametric equation of the straight line is



"x=t, y=-t, z=1"

Substitute



"t^2+(-t)^2-2(1)=C_1""t(-t)(1)=C_2"

Eliminate "t"



"2t^2-2=C_1""t^2=-C_2"

Then



"-2C_2-2=C_1"

Or



"C_1+2C_2+2=0"

Hence, the integral surface, which contains the straight line



"x^2+y^2-2z+2xyz+2=0"